if 1,log y x,log z y,-15log x z are in AP, then z 3 = x z -3 =y x = y -1 all of the above

if 1,logyx,logzy,-15logxz are in AP, then
  1. z= x
  2. z-3 =y
  3. x = y-1
  4. all of the above

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Log x, Logy ,Log z are in AP
2 Log y = logx + log z
log y^2 = log ( xz)
compare
y^2 = xz
…...................................